OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2. My function is f(x)=12x+6

OpenStudy (mathmale):

Hint: for this to happen, functions f(x) and h(x) must be inverses of each other. If one function is f(x)=12x+6, what is its inverse, h(x)?

OpenStudy (anonymous):

x-6/12=f(x) ???

OpenStudy (mathmale):

Have you found inverse functions before? if so, what do y ou remember of the procedure for doing so?

OpenStudy (anonymous):

I haven't done the inverse of functions before so I don't understand at all

OpenStudy (mathmale):

If you have no experience with inverse functions, then starting from scratch is a big order for you and a big order for me. Have you any course materials that show examples of finding inverse functions of given functions? Book? Notes? Online study materials?

OpenStudy (anonymous):

Yeah I have an online book..

OpenStudy (mathmale):

I did a quick Internet search for "finding inverse functions" and came up with the following: http://www.purplemath.com/modules/invrsfcn3.htm Have you taken the time to read the online book to learn something about inverse functions and how to find them?

OpenStudy (mathmale):

Please start with f(x) = 12x + 6. 1) replace f(x) with y. (please do this, and type in your result here.)

OpenStudy (anonymous):

y=x+6/12

OpenStudy (mathmale):

You are partially correct. Please be patient with me: Answer my questions. Please start with f(x) = 12x + 6. 1) replace f(x) with y. (please do this, and type in your result here.) Desired result: y=12x+6. 2) Interchange x and y. Result:

OpenStudy (anonymous):

y+6/12=x

OpenStudy (mathmale):

@uhhbianca...aren't you willing to answer the questions I'm asking you? To do so will take just a minute or two longer. Desired result: y=12x+6. 2) Interchange x and y. Result: ????? Your latest answer is partially correct. Why not stick with me and learn the whole procedure correctly?